# Advances in Radiance Field for Dynamic Scene: From Neural Field to Gaussian Field

### v1

@@ Abstract
Dynamic scene representation and reconstruction have undergone transformative advances in recent years, catalyzed by breakthroughs in neural radiance fields and 3D Gaussian splatting techniques. While initially developed for static environments, these methodologies have rapidly evolved to address the complexities inherent in 4D dynamic scenes through an expansive body of research. Coupled with innovations in differentiable volumetric rendering, these approaches have significantly enhanced the quality of motion representation and dynamic scene reconstruction, thereby garnering substantial attention from the computer vision and graphics communities. This survey presents a systematic analysis of over 200 papers focused on dynamic scene representation using radiance field, spanning the spectrum from implicit neural representations to explicit Gaussian primitives. We categorize and evaluate these works through multiple critical lenses: motion representation paradigms, reconstruction techniques for varied scene dynamics, auxiliary information integration strategies, and regularization approaches that ensure temporal consistency and physical plausibility. We organize diverse methodological approaches under a unified representational framework, concluding with a critical examination of persistent challenges and promising research directions. By providing this comprehensive overview, we aim to establish a definitive reference for researchers entering this rapidly evolving field while offering experienced practitioners a systematic understanding of both conceptual principles and practical frontiers in dynamic scene reconstruction. We maintain an active repository of literature and open-source implementations to complement this survey here.
@@
Dynamic Scene Representation, Neural Radiance Field, 3D Gaussian Splatting, Motion Representation, 4D Volume.
@@ 1 Introduction
Figure 1: Survey at A Glance. (a) Introduction and Foundation. We trace the evolution from static to dynamic scene representation, highlighting the challenges of jointly modeling motion, geometry, and appearance using radiance fields. (b) Motion Representation. We categorize motion patterns and their representation paradigms, examining how they enable complex motion modeling while addressing inherent limitations. (c) Scene Reconstruction. We analyze how motion representations enable scene reconstruction, discussing these methods within a unified framework while investigating how auxiliary information and regularization strategies constrain the learning of radiance fields. (d) Future Trends. We explore promising research directions and how dynamic scene reconstruction could benefit by aid of the rapid development of foundation models, and large language models. Partial photos courtesy of open-source datasets, and online resources.
@@
However, these early successes predominantly addressed static settings [1, 2, 4, 5, 6, 7], overlooking the inherent dynamics of real-world scenes. In practice, virtually every environment exhibits temporal evolution, whether from object movement, changing illumination, or evolving scene geometry. Recognizing this limitation, numerous recent approaches have extended static radiance field frameworks to handle dynamic scenes and accommodate complex temporal variations [8, 9, 10, 11, 12]. This rapidly expanding corpus of techniques underscores the necessity for a comprehensive survey synthesizing the state-of-the-art in dynamic scene representation.
The principal challenge in dynamic scene reconstruction lies in accurately modeling the temporal dimension–the motion field. Motion representation constitutes the cornerstone of dynamic scene reconstruction, where the precision of point correspondence across frames directly determines the quality of recovered dynamic content. To address this challenge, our survey begins by systematically examining the taxonomy of motion types and comprehensively reviewing strategies to represent these varies motion in 3D space. Recent advances have demonstrated that continuous and flexible formulations prove especially effective in faithfully representing complex motions without relying on oversimplified discretizations, as evidenced by innovations in neural scene flow fields, deformable radiance fields, and 4D neural volumes [13, 9, 14, 15, 11, 12].
@@
This survey aims to chart the evolutionary trajectory of dynamic scene representation, highlighting the substantial progress enabled by neural radiance fields and 3D Gaussian splatting while drawing attention to persistent challenges that require further investigation. Figure 1 provides a comprehensive overview of this survey’s structure and scope. By offering a systematically organized examination of recent innovations in motion representation and dynamic scene reconstruction, we seek to provide both newcomers and experienced researchers with valuable insights into emerging directions where dynamic scene modeling can evolve, ultimately facilitating increasingly realistic, interactive, and robust applications across computer vision, graphics, and related fields.
@@ 1.1 Roadmap
The field evolved further with the advent of 3DGS [2], which represent 3D scene with Gaussian primitives and rendered through efficient splatting techniques [29]. Replacing implicit neural representations with explicit Gaussian primitives, 3DGS demonstrate great potential for motion modeling. The approaches, that utilizing 3DGS to represent dynamic scenes, fall into several categories: methods that initialize Gaussians at time $t_{0}$ as a canonical space and warp them using time-dependent deformation fields [12, 43, 44]; techniques that track Gaussian primitive movements to represent dense motion fields [17, 45, 46]; approaches employing time-dependent functions to characterize varying Gaussian properties; and methods utilizing interpolated time-related features on factorized feature planes to predict dynamic Gaussian properties (position, scale, and rotation) [11, 47]. Recent advances have also applied 3D Gaussian fields to represent human avatars [31, 48, 49] and scenes with rigid objects [28, 50, 51, 52], enabling part-level animation or object-level manipulation.
@@ 1.2 Comparison to Related Surveys
Rapid progress in differentiable rendering and radiance field has led to numerous surveys in this field. The works of [53, 54, 55] provide foundational insight into differentiable rendering, while surveys in [56, 57, 58, 56, 59] document neural radiance field variants. More recently, surveys on 3D Gaussian splatting [60, 61, 62, 63] have emerged to capture developments in explicit representations.
However, a critical gap persists: existing surveys predominantly address static scene representation, with only peripheral coverage of dynamic scenes. Domain-specific surveys [64, 65, 66, 67, 68, 69] incorporate dynamic aspects, but focus on application-specific challenges rather than fundamental problems in dynamic scene representation. Even surveys explicitly addressing non-rigid reconstruction either take a broader view beyond neural fields [70, 71] or focus narrowly on specific scenarios[72, 73, 74]. Our survey distinguishes itself by providing a comprehensive analysis specifically dedicated to dynamic scene reconstruction using radiance fields, including both neural field and 3DGS. We uniquely bridge various representation paradigms within a unified framework, offering a new perspective on this rapidly evolving field.
@@ 2.1.2 Capture Setting
The methodology employed to obtain data substantially influences the quality of dynamic scene reconstruction, affecting both motion fidelity and temporal consistency. Capture settings can be categorized by the number of observation views (single-view, sparse-view, dense-view) or by the capture patterns in the scene (front-view, surrounding-view, or free-view). In this survey, we categorize capture settings into monocular, multi-view, and casual videos.
Monocular videos are captured using one camera, which could be stationary (single-view) or moving around (multi-view), but at any given time, only one view is available. As dynamic objects evolve along time, this setting suffers from severe ambiguities in geometry and motion, particularly for occluded regions, making it a challenging scenario for reconstruction.
Multi-view videos are captured by synchronized multiple cameras, yielding more comprehensive coverage and fewer occlusions. This configuration provides strong geometric constraints for reconstruction, enabling higher fidelity results but requiring sophisticated hardware setups and careful calibration, which limits its accessibility for everyday applications. Depending on the number of cameras, the system could be sparse or dense, while the position relative to the dynamic object could be front-facing or surrounding, each offering different trade-offs between coverage and complexity.
Casual videos are captured using hand-held devices without sophisticated setups, normally in free-view capture patterns and camera poses and parameters are often unavailable. These videos frequently suffer from low quality, motion blur, camera jitter, and complex background distractors that complicate dynamic scene analysis. While this setting offers maximum flexibility and accessibility, it must address not only the challenges of monocular capture (limited views at each timestep) but also the unstable camera setting, more noisy background, requiring robust algorithms for camera tracking, foreground-background separation, and temporal coherence maintenance.
Figure 3: Relative motion speed between camera and objects determines the critical factor in effective scene capture.
2.1.3 Discussion
Dynamic scene reconstruction requires tracking points across video frames, where 2D pixel movement represents a combination of both object and camera motion. The type of sensor setup and capture strategy significantly impacts the system’s ability to accurately estimate true 3D motion, ultimately determining the upper bound of reconstruction outcomes. For static scenes, a single moving camera can provide information equivalent to multiple cameras by capturing different viewpoints over time. However, when objects themselves move, the relationship between camera motion and object motion becomes crucial [75]. This relationship creates different levels of ambiguity and reconstruction difficulty, categorized as: (1) strict monocular, where the camera moves much slower than objects, creating occlusion challenges, (2) effective multi-view, where camera and object speeds are comparable, allowing the camera to “follow" and maintain visibility of key points, and (3) strict multi-view, where multiple synchronized cameras capture the scene simultaneously from different views. When camera movement closely matches object movement, even single-camera setups can track features effectively enough for reasonable reconstruction, as illustrated in Fig. 3. Importantly, the critical factor is not the absolute motion speed of camera or object, but rather their relative speed ratio.
While strict multi-view camera setups theoretically provide the most complete information for high-quality reconstruction, they introduce practical challenges in synchronization, data management, and deployment complexity. Strict monocular approaches, despite capturing less complete information about the scene, offer significantly simpler solutions for real-world applications. The trade-off between capture complexity and reconstruction quality continues to shift as algorithms improve in handling limited input data, with effective multi-view approaches representing a promising middle ground, making simpler camera setups increasingly viable for many practical uses.
@@ 2.2 Volumetric Radiance Field
Volumetric radiance fields have emerged as powerful representations for novel view synthesis, offering compelling alternatives to traditional mesh-based or voxel-based scene representations. These approaches encode both geometry and appearance information, enabling photorealistic rendering from arbitrary viewpoints while being directly optimized from multi-view imagery without requiring explicit 3D reconstruction. In this survey, we mainly focuses on two prominent implementations: Neural Radiance Fields (NeRF) and 3D Gaussian Splatting.
@@ 2.2.1 Neural Radiance Field
$\mathcal{F}_{\theta}(\gamma(\mathbf{x}),\mathbf{d})\longmapsto(\mathbf{c},\sigma),$ (1)
where $\mathbf{x}=(x,y,z)$ denotes spatial coordinates, $\gamma(\cdot)$ is a positional encoding function that maps inputs to higher-dimensional space to capture high-frequency details, and $\mathbf{d}=(d_{x},d_{y},d_{z})$ represents viewing direction as a normalized unit vector from the camera’s optical center to a pixel position. The network outputs RGB color $\mathbf{c}=(r,g,b)$ and volume density $\sigma$, where density represent view-independent geometry, while color varies with viewing direction to model view-dependent effects such as specular highlights.
Volume Rendering. The rendering process in NeRF employs ray tracing principles [24], integrating color and density values along camera rays $\mathbf{r}(t)=\mathbf{o}+t\mathbf{d}$ to produce the final pixel color. This integration is expressed as:
[…]
@@
An alternative approach to volumetric scene representation is offered by 3D Gaussian Splatting (3DGS) [2], which employs explicit, learnable primitives rather than implicit neural networks. This method represents scenes as collections of anisotropic 3D Gaussians $\mathcal{G}$, each parameterized by its position $\boldsymbol{\mu}\in\mathbb{R}^{3}$, covariance matrix $\boldsymbol{\Sigma}\in\mathbb{R}^{3\times 3}$, opacity $o\in[0,1]$, and color attributes $\mathbf{c}$. The covariance matrix, defining the Gaussian’s shape and orientation, is constructed from a scaling factor $\boldsymbol{S}\in\mathbb{R}^{3}$ and rotation matrix $\boldsymbol{R}\in\mathbb{R}^{3\times 3}$ as $\boldsymbol{\Sigma}=\boldsymbol{R}\boldsymbol{S}\boldsymbol{S}^{\top}\boldsymbol{R}^{\top}$. The color attributes are typically represented by spherical harmonics (SH) coefficients to model view-dependent appearance effects. All these properties are learnable parameters that are optimized through gradient descent to align with multi-view imagery, allowing the model to accurately represent both geometry and appearance. The explicit nature of this representation offers advantages in terms of interpretability and direct manipulation, though complex scenes may require large numbers of Gaussians, potentially increasing memory requirements.
Volume Splatting. Unlike NeRF’s ray marching approach, Gaussian Splatting employs a tile-based rasterization pipeline for efficient rendering [29]. The process involves projecting 3D Gaussian primitives onto the 2D image plane, a technique commonly referred to as "splatting." When projecting each 3D Gaussian, its center is transformed to the 2D image space as $\boldsymbol{\mu}^{2D}=\boldsymbol{J}\boldsymbol{W}\boldsymbol{\mu}$, and its covariance matrix is similarly transformed to $\boldsymbol{\Sigma}^{2D}=\boldsymbol{J}\boldsymbol{W}\boldsymbol{\Sigma}\boldsymbol{W}^{\top}\boldsymbol{J}^{\top}$, where $\boldsymbol{W}$ represents the viewing transformation matrix that maps from world to camera coordinates, and $\boldsymbol{J}$ is the Jacobian of the affine approximation of the projective transformation.
@@
where $\mathcal{N}$ represents the set of Gaussians contributing to the pixel, $\mathbf{c}_{i}$ is the color of the $i$-th Gaussian, and $\alpha_{i}^{2D}$ is the alpha value derived from the Gaussian’s learned opacity $o_{i}$ and its 2D projection. This alpha value is calculated as:
$\alpha_{i}^{2D}=o_{i}\exp\left(-\frac{1}{2}(\mathbf{x}^{2D}-\boldsymbol{\mu}_{i}^{2D})^{\top}{(\boldsymbol{\Sigma}_{i}^{2D}})^{-1}(\mathbf{x}^{2D}-\boldsymbol{\mu}_{i}^{2D})\right),$ (5)
where $\mathbf{x^{2D}}$ represents the 2D coordinates of the pixel being rendered.
This splatting-based approach offers significant computational advantages over ray marching, enabling real-time rendering while maintaining high visual fidelity. The explicit nature of the Gaussian primitives also facilitates adaptive density control during optimization, allowing the model to concentrate more Gaussians in complex regions of the scene while using fewer in simpler areas. Moreover, the differentiable nature of the entire pipeline enables end-to-end optimization using only multi-view images as supervision, without requiring explicit 3D ground truth.
2.2.3 Discussion
Neural Radiance Fields and 3D Gaussian Splatting represent complementary approaches to volumetric scene representation, each with distinct tradeoffs. NeRF’s implicit continuous representation excels at capturing fine details and complex view-dependent effects through compact network parameters, but requires computationally expensive ray marching that precludes real-time applications. In contrast, 3D Gaussian Splatting’s explicit primitive-based representation enables real-time rendering through efficient rasterization, though complex scenes may demand large numbers of Gaussians that increase memory requirements. The choice between these approaches thus depends on specific application priorities—whether visual fidelity, computational efficiency, or memory constraints takes precedence.
These methods fundamentally differ in their rendering techniques: ray tracing in NeRF provides continuous sampling that accurately captures complex optical phenomena, while splatting in Gaussian representations serves as an approximation of physically-based rendering, sacrificing some fidelity in representing sophisticated light transport effects such as reflections and shadows. This complementarity has motivated recent exploration of hybrid techniques that introduce ray tracing capabilities into Gaussian-based representations to better capture secondary lighting effects while maintaining computational efficiency [76, 77]. As the field advances, developing representations that simultaneously achieve efficiency, fidelity, and real-time performance remains an active frontier in neural rendering, with each approach continuing to inform and enhance the other.
@@ 3 Dynamic Motion Representation
Accurately modeling dynamic motion forms a critical foundation for scene reconstruction, understanding, and analysis. Real-world environments exhibit diverse motion patterns that can be categorized hierarchically from specific to general types. We classify these patterns into rigid motion, articulated motion, general non-rigid motion, and hybrid motion (which combines multiple patterns). Although articulated motion represents a specific type of non-rigid movement, we address it separately due to its distinct representation approaches [78]. Throughout this survey, we use the term “non-rigid” to describe more general deformation patterns beyond articulated movement.
@@
$\mathbf{x}_{t}\;=\;\mathcal{T}_{\theta}\bigl{(}\mathbf{x}_{t-1}\,\mid\,\pi\bigr{)},$ (6)
where $\mathcal{T}_{\theta}(\cdot)$ is a transformation function parameterized by $\theta$ and conditioned on temporal context $\pi$ (e.g., time $t$, frame index $i$, embeddings $\ell_{t}$, or motion-specific parameters). The precise form of $\mathcal{T}_{\theta}(\cdot)$ depends on the underlying motion pattern and representation method, typically involving different assumptions and formulations.
Figure 4: A 2D illustration of various motion types with representative examples.

### v2

@@ Abstract
Dynamic scene representation and reconstruction have undergone transformative advances in recent years, catalyzed by breakthroughs in neural radiance fields and 3D Gaussian splatting techniques. While initially developed for static environments, these methodologies have rapidly evolved to address the complexities inherent in 4D dynamic scenes through an expansive body of research. Coupled with innovations in differentiable volumetric rendering, these approaches have significantly enhanced the quality of motion representation and dynamic scene reconstruction, thereby garnering substantial attention from the computer vision and graphics communities. This survey presents a systematic analysis of over 200 papers focused on dynamic scene representation using radiance field, spanning the spectrum from implicit neural representations to explicit Gaussian primitives. We categorize and evaluate these works through multiple critical lenses: motion representation paradigms, reconstruction techniques for varied scene dynamics, auxiliary information integration strategies, and regularization approaches that ensure temporal consistency and physical plausibility. We organize diverse methodological approaches under a unified representational framework, concluding with a critical examination of persistent challenges and promising research directions. By providing this comprehensive overview, we aim to establish a definitive reference for researchers entering this rapidly evolving field while offering experienced practitioners a systematic understanding of both conceptual principles and practical frontiers in dynamic scene reconstruction. We maintain an active repository of literature and open-source implementations to complement this survey at Awesome-DynRF.
@@
Motion Representation, Dynamic Scenes, Neural Radiance Field, 3D Gaussian Splatting.
@@ 1 Introduction
Figure 1: Survey at A Glance. (a) Introduction and Foundation. We trace the evolution from static to dynamic scene representation, highlighting the challenges of jointly modeling motion, geometry, and appearance using radiance fields. (b) Motion Representation. We categorize motion patterns and their representation paradigms, examining how they enable complex motion modeling while addressing inherent limitations. (c) Scene Reconstruction. We analyze how motion representations enable scene reconstruction, discussing these methods within a unified framework while investigating how auxiliary information and regularization strategies constrain the learning of radiance fields. (d) Future Trends. We explore promising research directions and how dynamic scene reconstruction could benefit by aid of the rapid development of foundation models and large language models.
@@
However, these early successes predominantly addressed static settings [1, 2, 4, 5, 6, 7], overlooking the inherent dynamics of real-world scenes. In practice, virtually every environment exhibits temporal evolution, whether from object movement, changing illumination, or evolving scene geometry. Recognizing this limitation, numerous recent approaches have extended static radiance field frameworks to handle dynamic scenes and accommodate complex temporal variations [8, 9, 10, 11, 12]. This rapidly expanding corpus of techniques underscores the necessity for a comprehensive survey summarizing the state-of-the-art in dynamic scene representation and reconstruction.
The principal challenge in dynamic scene reconstruction lies in accurately modeling the temporal dimension–the motion field. Motion representation constitutes the cornerstone of dynamic scene reconstruction, where the precision of point correspondence across frames directly determines the quality of recovered dynamic content. To address this challenge, our survey begins by systematically examining the taxonomy of motion types and comprehensively reviewing strategies to represent these various motions in 3D space. Recent advances have demonstrated that continuous and flexible formulations prove especially effective in faithfully representing complex motions without relying on oversimplified discretizations, as evidenced by innovations in neural scene flow fields, deformable radiance fields, and 4D neural volumes [13, 9, 14, 15, 11, 12].
@@
This survey aims to chart the evolutionary trajectory of dynamic scene representation, highlighting the substantial progress enabled by neural radiance fields and 3D Gaussian splatting while drawing attention to persistent challenges that require further investigation. Fig 1 provides a comprehensive overview of this survey’s structure and scope. By offering a systematically organized examination of recent innovations in motion representation and dynamic scene reconstruction, we seek to provide both newcomers and experienced researchers with valuable insights into emerging directions where dynamic scene modeling can evolve, ultimately facilitating increasingly realistic, interactive, and robust applications across computer vision, graphics, and related fields.
@@ 1.1 Roadmap
The field evolved further with the advent of 3DGS [2], which represents 3D scene with Gaussian primitives and is rendered through efficient splatting techniques [29]. Replacing implicit neural representations with explicit Gaussian primitives, 3DGS demonstrates great potential for motion modeling. The approaches, that utilizing 3DGS to represent dynamic scenes, fall into several categories: methods that initialize Gaussians at time $t_{0}$ as a canonical space and warp them using time-dependent deformation fields [12, 43, 44]; techniques that track Gaussian primitive movements to represent dense motion fields [17, 45, 46]; approaches employing time-dependent functions to characterize varying Gaussian properties; and methods utilizing interpolated time-related features on factorized feature planes to predict dynamic Gaussian properties (position, scale, and rotation) [11, 47]. Recent advances have also applied 3D Gaussian fields to represent human avatars [31, 48, 49] and scenes with rigid objects [28, 50, 51, 52], enabling part-level animation or object-level manipulation.
@@ 1.2 Comparison to Related Surveys
Rapid progress in differentiable rendering and radiance field has led to numerous surveys in this field. The works of [53] provide foundational insight into differentiable rendering, while surveys in [54, 55, 54] document neural radiance field variants. More recent surveys on 3DGS [56, 57, 58] have emerged to capture developments in explicit representations.
However, a critical gap persists: existing surveys predominantly address static scene representation, with only peripheral coverage of dynamic scenes. Domain-specific surveys [59, 60, 61, 62, 63] incorporate dynamic aspects, but focus on application-specific challenges rather than fundamental problems in dynamic scene representation. Even surveys explicitly addressing non-rigid reconstruction either take a broader view beyond neural fields [64, 65] or focus narrowly on specific scenarios[66, 67]. Our survey distinguishes itself by providing a comprehensive analysis specifically dedicated to dynamic scene reconstruction using radiance fields, including NeRF and 3DGS. We uniquely bridge various representation paradigms within a unified framework, offering a new perspective on this rapidly evolving field.
@@ 2.1.2 Capture Setting
We categorize scene capture settings into three distinct classes: monocular capture which encompasses both strict monocular with stationary cameras and effective multi-view when camera motion is comparable to object speed, multi-view capture that employs multiple synchronized cameras that simultaneously observe the scene from different angles, providing comprehensive geometric constraints, and casual video capture, footage obtained from handheld devices in unconstrained environments without professional setups. The key distinction between these approaches lies in their spatiotemporal sampling characteristics and the relative motion between camera and scene objects [68]. While multi-view setups offer superior reconstruction fidelity through comprehensive spatial coverage with minimal occlusions, monocular methods can achieve reasonable results when relative camera-object motion is appropriately balanced, and casual video approaches trade reconstruction quality for accessibility and flexibility in everyday scenarios.
@@ 2.2.1 Neural Radiance Field
$\mathcal{F}_{\theta}(\mathbf{x},\mathbf{d})\longmapsto(\mathbf{c},\sigma),$ (1)
where $\mathbf{x}=(x,y,z)$ denotes spatial coordinates, and $\mathbf{d}$ represents viewing directions as normalized unit vectors from the camera’s optical center to pixel positions. The network outputs RGB color $\mathbf{c}=(r,g,b)$ and volume density $\sigma$, where density represents view-independent geometry, while color varies with viewing direction to model view-dependent effects such as specular highlights.
Volume Rendering. The rendering process in NeRF employs ray tracing principles [24], integrating color and density values along camera rays $\mathbf{r}(t)=\mathbf{o}+t\mathbf{d}$ to produce the final pixel color. This integration is expressed in discrete form:
@@
3DGS [2] provides an alternative to employs explicit, learnable primitives rather than implicit neural networks to represent the radiance field. This method represents scenes as collections of anisotropic 3D Gaussians $\mathcal{G}$, each parameterized by its position $\mathbf{\mu}\in\mathbb{R}^{3}$, covariance matrix $\mathbf{\Sigma}\in\mathbb{R}^{3\times 3}$, opacity $o\in[0,1]$, and color attributes $\mathbf{c}$. The covariance matrix, defining the Gaussian’s shape and orientation, is constructed from a scaling factor $\mathbf{S}\in\mathbb{R}^{3}$ and rotation matrix $\mathbf{R}\in\mathbb{R}^{3\times 3}$ as $\mathbf{\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{\top}\mathbf{R}^{\top}$. The color attributes are typically represented by spherical harmonics (SH) coefficients to model view-dependent appearance effects. All these properties are learnable parameters that are optimized through gradient descent to align with observations.
Volume Splatting. Unlike NeRF’s ray marching approach, Gaussian Splatting employs a tile-based rasterization pipeline for efficient rendering [29]. The process involves projecting 3D Gaussian primitives onto the 2D image plane, a technique commonly referred to as "splatting." When projecting each 3D Gaussian, its center is transformed to the 2D image space as $\mathbf{\mu}^{2D}=\mathbf{J}\mathbf{W}\mathbf{\mu}$, and its covariance matrix is similarly transformed to $\mathbf{\Sigma}^{2D}=\mathbf{J}\mathbf{W}\mathbf{\Sigma}\mathbf{W}^{\top}\mathbf{J}^{\top}$, where $\mathbf{W}$ represents the viewing transformation matrix that maps from world to camera coordinates, and $\mathbf{J}$ is Jacobian of the affine approximation of projective transformation.
@@
where $\mathcal{N}$ represents the set of Gaussians contributing to the pixel, $\mathbf{c}_{i}$ is the color of the $i$-th Gaussian, and $\alpha_{i}^{2D}$ is the 2D projected alpha value of each gaussian primitive.
@@ 3 Dynamic Motion Representation
Accurately modeling dynamic motion forms a critical foundation for scene reconstruction, understanding, and analysis. Real-world environments exhibit diverse motion patterns that can be categorized hierarchically from specific to general types. We classify these patterns into rigid motion, articulated motion, general non-rigid motion, and hybrid motion, which combines multiple patterns, as illustrated in Fig. 3. Although articulated motion represents a specific type of non-rigid movement, we address it separately due to its distinct representation approaches [69]. Throughout this survey, we use the term ”non-rigid” to describe more general deformation patterns beyond articulated movement.
@@
$\mathbf{x}_{t}\;=\;\mathcal{T}_{\theta}\bigl{(}\mathbf{x}_{t-1};\pi\bigr{(}t)),$ (4)
where $\mathcal{T}_{\theta}(\cdot)$ is a transformation function parameterized by $\theta$ and conditioned on temporal context $\pi$ (e.g., time $t$, frame index $i$, latent code $\ell_{t}$, or motion-specific parameters). The precise form of $\mathcal{T}_{\theta}(\cdot)$ depends on the underlying motion pattern and representation method, typically involving different assumptions and formulations.
Figure 3: A 2D illustration of various motion types.